OpenStudy (manny_9):

If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 6. -30 -32 -34 -36

OpenStudy (manny_9):

@Directrix

Directrix (directrix):

y = k*x where k is the constant of proportionality. y = k*x 15 = k*(-3) k = ? What is k = ? @manny_9

OpenStudy (manny_9):

-5?

Directrix (directrix):

Yes.

Directrix (directrix):

Build the equation y = -5 * 6 where k = -5 and x = 6. y = ?

OpenStudy (manny_9):

-30

Directrix (directrix):

That is what I got. Correct.

OpenStudy (manny_9):

Thanks can u check this http://prntscr.com/cswppv

Directrix (directrix):

Direct Variation is correct. Inverse variation has a curved graph.

OpenStudy (manny_9):

Thanks

Directrix (directrix):

You're welcome.